OpenStudy (anonymous):

Guys need help please... Solve for x: √(5x-4)=x/2+2

OpenStudy (anonymous):

First Raise both sides to the power of two: \[5x-4=(x/2+2)^2\] expand terms on right side: \[5x-4=(x^2/4)+2x+4\] move everything to left side: \[(-x^2/4+3x-8)=0\] the left hand side factors into: -1/4((x-8)(x-4))=0 multiply both sides by -4 (x-4)(x-8)=0 So the two answers are x =4 and x=8.

OpenStudy (anonymous):

Note that you'll have to restrict the domain of the function and check if the answer are in the domain after you solve it.

OpenStudy (anonymous):

Both of those are in the domain.

OpenStudy (anonymous):

:) Then there's nothing wrong. Sorry, it's a habit I got from my strict math teacher.

OpenStudy (anonymous):

Oh no need to apologize, I went back and checked it after you said that.

OpenStudy (anonymous):

lol Maths is fun.

OpenStudy (anonymous):

Thank you!! ((:

OpenStudy (anonymous):

Can I have a medal?

OpenStudy (anonymous):

sure! but how can i give it to you?